# Analytical exciton energies in monolayer transition-metal dichalcogenides

### v2

@@ Abstract
We derive an analytical expression for $s$-state exciton energies in monolayer transition-metal dichalcogenides (TMDCs): $E_{\text{ns}}=-{\text{Ry}}^{*}\times P_{n}/{(n-1/2+0.479\,r^{*}_{0}/\kappa)^{2}}$, $n=1,2,...$, where $r^{*}_{0}$ and $\kappa$ are the dimensionless screening length and dielectric constant of the surrounding medium; $\text{Ry}^{*}$ is an effective Rydberg energy scaled by the dielectric constant and exciton reduced mass; $P_{n}(r^{*}_{0}/\kappa)$ is a function of variables $n$ and $r^{*}_{0}/\kappa$. Its values are around 1.0 so we consider it a term that corrects the Rydberg energy. Despite the simple form, the suggested formula gives exciton energies with high precision compared to the exact numerical solutions that accurately describe recent experimental data for a large class of TMDC materials, including WSe2, WS2, MoSe2, MoS2, and MoTe2. To achieve these results, we have developed a so-called regulated perturbation theory by combining the conventional perturbation method with several elements of the Feranchuk-Komarov operator method, including the Levi-Civita transformation, the algebraic calculation technique via the annihilation and creation operators, and the introduction of a free parameter to optimize the convergence rate of the perturbation series. This universal form of exciton energies could be helpful in various physical analyses, including retrieval of the material parameters such as reduced exciton mass and screening length from the available measured exciton energies.
@@
Introduction – In the last decade, exciton energies in different monolayer materials have been measured and studied theoretically simultaneously [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Pioneer studies [1, 2] indicate that unlike 3D bulk semiconductors, electrons and holes in 2D atomically thin films such as monolayer transition-metal dichalcogenides (TMDCs) couple to each other by Rytova-Keldysh interaction [17, 18, 19, 20] when forming neutral excitons. As this interaction behaves as logarithm potential $\ln(r)$ in the short-range regime and becomes a conventional Coulomb interaction $1/r$ in the long-range regime [18, 20], the exciton energies of low-lying states ($n<5$) do not obey the two-dimensional hydrogen atom’s Ryberg series. Thus, an accurate theory of exciton binding energies in 2D semiconductors is necessary to benchmark experiments.
Recently, different approaches have been developed to determine exciton binding energies when the reduced mass, screening length, and dielectric constant of the surrounding medium are known [12, 13, 14, 15]. In particular, our packages [13, 14, 15] based on the Feranchuk-Komarov operator method [21, 22] perform excellently in calculating exciton binding energies in monolayer TMDCs, even in a high magnetic field. On the other hand, some theoretical efforts were made to obtain analytical formulas for exciton binding energies in 2D semiconductors [5, 6, 7]. These formulas arise from modeling the electron-hole interaction potential in 2D semiconductors to make the associated Wannier Schrödinger equation analytically solvable. This tactic leads to easy-to-use formulas generalized from the Rydberg series with acceptable accuracy. However, of course, the accuracy of these formulas is limited by the accuracy of their modified electron-hole potentials. Therefore, a comprehensive theory is needed to describe exciton energies analytically.
In this Letter, we provide highly accurate analytical expressions for $s$-state exciton energies based on a novel approach, the regulated perturbation theory (RPT). The obtained formulas for exciton energies, simple but universal for a wide range of TMDC monolayers, are precise compared with the exact numerical solutions and describe the currently available experimental data well. This achievement is granted by the advances of the method. Unlike conventional perturbation theory, the RPT approach includes the Levi-Civita transformation, which allows for the algebraic calculation technique and introduces a free parameter to control the convergence rate of the perturbation correction series. Moreover, it does not need to modify the electron-hole potential and precisely calculate all matrix elements concerning the harmonic oscillator basis set.
@@
where $\alpha=r_{0}/\kappa a_{0}^{*}$ is a dimensionless parameter dependent on the screening length $r_{0}$.
@@
is excluded from the Coulomb singularity, making solving it more convenient. Notably, it allows the solving process to use the harmonic oscillator basis set $|n,m(\omega)\rangle$ with running indices $n=0,1,2,\cdots$ and $m=0,\pm 1,\pm 2,\cdots,\pm n$ (the principal and magnetic quantum numbers). The harmonic frequency $\omega$ can be used as a free parameter to regulate the convergence rate of the perturbation series so that the desired precision is granted. Moreover, the equation in $(u,v)$ space and the harmonic basis set can be presented via the creation and annihilation operators: $\hat{a}^{+},\hat{a}$ and $\hat{b}^{+},\hat{b}$. All matrix elements necessary for solving Eq. (3), such as $h_{k}=\langle n|\hat{H}|n\rangle$, $r_{k}=\langle n|\hat{R}|n\rangle$, $v_{jk}=\langle j|\hat{H}|k\rangle$, and $v^{\text{\tiny{\it R}}}_{jk}=\langle j|\hat{R}|k\rangle$ $(j\neq k)$, can then be calculated algebraically using the commutation relations $[\hat{a},\hat{a}^{+}]=1$ and $[\hat{b},\hat{b}^{+}]=1$. Here, we consider only $s$-states where $m=0$, so the basis set can be shortly written as $|n\rangle\equiv|n,m=0\rangle$. This algebraic approach obtained analytical forms of these matrix elements in Ref. [15]. We provide them in Appendix A for use in the present study.
@@
Different from Ref. [15], where Eq. (3) was treated numerically with highly accurate exciton energies, we solve this equation analytically by the perturbation method, expanding wave functions and energies by the perturbation parameter $\beta$ as $\psi=\psi^{(0)}+\beta\Delta\psi^{(1)}+\beta^{2}\Delta\psi^{(2)}+\cdots$, $E=E^{(0)}+\beta\Delta E^{(1)}+\beta^{2}\Delta E^{(2)}+\cdots$. Plugging these expansions into Eq. (3), where the operators are separated as $\hat{H}=\hat{H}_{0}+\beta\,\hat{V}$ and $\hat{R}=\hat{R}_{0}+\beta\hat{V}_{R}$, we obtain formulas for the wave function $\psi^{(0)}$ and energy $E^{(0)}$ in the zero-order approximation and their corrections $\Delta\psi^{(s)}$, $\Delta E^{(s)}$ in any approximation order $s=1,2,3,\cdots$. The present study is interested in exciton energies only to the third-order approximation, whose formulas are
$E^{(0)}_{n}$ $=$ $\varepsilon_{n}=\frac{h_{n}}{r_{n}},\quad\quad\Delta E_{n}^{(1)}=0,$
$\Delta E_{n}^{(2)}$ $=$ $-\sum_{\begin{smallmatrix}j=0\\ j\neq n\end{smallmatrix}}^{+\infty}\dfrac{(v_{nj}-\varepsilon_{n}v^{\text{\tiny{\it R}}}_{nj})(v_{jn}-\varepsilon_{n}v^{\text{\tiny{\it R}}}_{jn})}{(h_{j}-\varepsilon_{n}r_{j}){r_{n}}},$ (4)
$\Delta E^{(3)}_{n}$ $=$ $\sum_{\begin{smallmatrix}k=0\\ k\neq n\end{smallmatrix}}^{+\infty}{\sum_{\begin{smallmatrix}j=0\\ j\neq k,n\end{smallmatrix}}^{+\infty}\dfrac{(v_{nk}-\varepsilon_{n}v^{\text{\tiny{\it R}}}_{nk})P_{kj}(v_{jn}-\varepsilon_{n}v^{\text{\tiny{\it R}}}_{jn})}{(h_{k}-\varepsilon_{n}r_{k})(h_{j}-\varepsilon_{n}r_{j}){r_{n}}}}.$
Here, we use the notation $P_{kj}=v_{kj}-\varepsilon_{n}v^{\text{\tiny{\it R}}}_{kj}$ for short; energies in the $s$-order approximation are calculated by equation $E^{(s)}=E^{(0)}+\Delta E^{(1)}+\Delta E^{(2)}+\cdots+\Delta E^{(s)}$.
@@
Table 1 shows exciton energies in monolayer WSe2 in the zero-, second-, and third-order approximations for $1s$ to $5s$ states, calculated by formulas (Analytical exciton energies in monolayer transition-metal dichalcogenides) and compared with the numerically exact solutions [15]. The free parameter $\omega$ is chosen by the variational condition $\partial E^{(0)}/\partial\omega=0$ for these energies. As shown in the Table, the energy precision is very high, within 1.0 meV for the second-order approximation, while less than 0.2 meV for the third-order approximation. We note that the current limit of the experimental measurement is about 1.0 meV.
@@
Table 1: Exciton energies (meV) in monolayer WSe2 encapsulated by hBN slabs with screening length $r_{0}=4.21$ nm, exciton reduced mass $\mu=0.19\,m_{e}$, and dielectric constant $\kappa=4.34$, calculated by Eqs. (Analytical exciton energies in monolayer transition-metal dichalcogenides) and compared to the exact numerical solutions [15].
$E^{(0)}$ $E^{(2)}$ $E^{(3)}$ $E_{num.}$ [15]
@@
We have also verified the accuracy of formulas (Analytical exciton energies in monolayer transition-metal dichalcogenides) numerically for some other TMDC monolayers: WS2, MoSe2, and MoS2. It shows the same high precision, less than $0.5$ meV, in the third-order approximation, which is not provided here but is available in Supplementary [24].
Analytical exciton energies – As shown above, formulas (Analytical exciton energies in monolayer transition-metal dichalcogenides) for the corrections of up to the third-order approximation give extremely accurate exciton energies with precision within 0.5 meV compared to the numerically exact solutions. Even for the energies in the second-order approximation, the precision is about 1.0 meV, a limit for experimental measurement. Therefore, we consider them for analytical exciton energies. Indeed, by substituting the explicit forms of matrix elements $h_{k}$, $r_{k}$, $v_{jk}$, and $v^{\text{\tiny{\it R}}}_{jk}$ from Appendix A into Eqs. (Analytical exciton energies in monolayer transition-metal dichalcogenides), we get particular exciton energies in the zero-order approximation as
$E^{(0)}_{n}=\frac{\kappa^{2}}{4{r_{0}^{*}}^{2}}\frac{1}{\lambda^{2}}-\frac{2\kappa}{r_{0}^{*}}f_{n}(\lambda).$ (5)
Here, $r_{0}^{*}=r_{0}/a_{0}^{*}$ is defined as a dimensionless screening length; $\lambda$ is a variational parameter defined from the free parameter $\omega$ by $\lambda=\kappa/\omega r_{0}^{*}$; $f_{n}(\lambda)$ is a function dependent on $\lambda$ only, whose explicit form is given in Appendix A. For convenience, from now on, the principle quantum number $n$ is counted from 1 (instead of zero as in the previous calculations) to 5, corresponding to the states from $1s$ to $5s$.
@@
$\frac{\partial E^{(0)}_{n}}{\partial\lambda}=0.$ (6)
We must solve this equation to get $\lambda$ and plug it into Eq. (5) to get the needed analytical energies. We demonstrate the solving process for $ns$ states ($n\leq 5$) in Supplementary [24] and provide only the main idea and final results in the main text.
Because of the complexity of function $f_{n}(\lambda)$, analytically solving Eq. (6) is not trivial, requiring the Taylor expansion of this function. From another side, numerically solving this equation for specific cases of monolayers WSe2, WS2, MoSe2, and MoS2 encapsulated by hBN slabs with the material parameters given in Table 3, we get particular solutions for $1s$ state: $\lambda=0.803,0.863,0.582,$ and $0.750$, corresponding to each material. These results suggest we expand function $f_{1}(\lambda)$ around the average value $\lambda_{0}=0.75$. This expansion leads Eq. (6) to a polynomial equation concerning an unknown $\delta\lambda=\lambda-0.75$, making it easy to get approximate analytical solutions consistent with the numerical ones. Analogically, we get $\lambda$ for other $ns$ states by expanding functions $f_{n}(\lambda)$ around the points $\lambda_{0}=(n+1/2)/2$ and then solving Eq. (6). We combine all the solutions in one generalized formula as
$\lambda$ $=$ $\frac{1}{2}\left(n+0.5\right)+\delta\lambda,$ (7)
with
$\delta\lambda$ $=$ $\left(0.47\,{\kappa}/{r_{0}^{*}}-0.525\right)n\,.$ (8)
We note that the actual solutions of Eq. (6) are not precisely as given by formula (8); however, a slight modification has been made to get a general formula for all states $ns$. This tactic is reasonable because exact energies should not depend on the free parameter, and its modification around the extremum point does not change the solutions much. Reference [22] provides a detailed discussion of how to choose such a free parameter.
@@
$E_{\text{ns}}$ $=$ $-\frac{{\text{Ry}}^{*}\times P_{n}}{(n-0.5+0.479\,r_{0}^{*}/\kappa)^{2}}.$ (9)
@@
with $\delta=(0.47\,{\kappa}/{r_{0}^{*}}-0.525)$ and coefficients $a_{n}$ and $b_{n}$ defined as
$a_{n}$ $=$ $4.3/n-{6.6}/{n^{2}}+{3.2}/{n^{3}},$
$b_{n}$ $=$ $5.3/n-{7.6}/{n^{2}}+{4.0}/{n^{3}}\,.$ (11)
$P_{n}$ is a function of variable $r_{0}^{*}/\kappa$ only with values around 1.0 and $P_{n}\rightarrow 1$ when $n\rightarrow+\infty$, so we can interpret it as a term correcting the effective Rydberg energy Ry.
@@
Validation with the available experimental data – To obtain analytical energies (9), we expand the equation around specific points $\lambda_{0}=(n+0.5)/2$ with the condition $\lambda\gg\delta\lambda$, which defines the working range of $r_{0}^{*}/\kappa$. From Eq. (8) for $\delta\lambda$, we get the following boundary for our theory
$0.89>r_{0}^{*}/\kappa\gg 0.46\quad\text{or}\quad 0.89<r_{0}^{*}/\kappa\ll 18.8$ (12)
@@
Furthermore, we verify formula (9) by comparing it with experimental data for monolayers WSe2 [9], WS2, MoSe2, MoS2, and MoTe2 [10]. The comparison in Fig. S-1 demonstrates a high consistency between the experiments and theory. Experimental data of Liu et al. [8] for monolayer WSe2 also support our theory: $\Delta E_{2s-1s}=131$(130.3) meV, $\Delta E_{3s-2s}=21$ (21.7) meV, $\Delta E_{4s-2s}=30$ (29.0) meV, where the values in parenthesis are theoretical from Eq. (9).
@@
Figure 1: Exciton energies for the $ns$ states in monolayers WSe2, WS2, MoSe2, MoS2, and MoTe2 calculated by formula (9) (horizontal axis) compared with the experimental data [10, 9] (vertical axis). Material parameters used for calculations are provided in Table 3.
@@
Conclusion – A novel approach based on the perturbation theory has been developed, including some elements of the FK operator method, to achieve $s$-state exciton energies in monolayer transition-metal dichalcogenides. The third-order approximation gives almost exact solutions, while the second-order approximation of the perturbation theory gives exciton energies with high precision, within 1.0 meV compared to the exact numerical solutions. This accuracy suggests we use them for analytical exciton energies. Although simple and universal for $s$-states, the constructed analytical formula describes experimental exciton energies for a wide range of monolayer TMDCs. This general approach can be applied to other kinds of 2D semiconduction materials.
Acknowledgments
@@
This work was funded by Vietnam Ministry of Education and Training under grant number B2022-SPS-09-VL and carried out by the high-performance cluster at Ho Chi Minh City University of Education, Vietnam.
A Material parameters and matrix elements
In this study, we use material parameters in Table 3
[…]
$\kappa$ $\mu$ ($m_{e}$) $r_{0}$ (nm) Ry (meV) $a^{*}_{0}$ (nm) Refs.
WSe2 4.34 0.19 4.21 137.24 1.209 [15]
WS2 4.16 0.175 3.76 137.59 1.258 [15]
MoSe2 4.40 0.35 3.90 245.97 0.665 [10]
MoS2 4.45 0.275 3.40 188.94 0.856 [10]
MoTe2 4.30 0.37 6.10 272.26 0.615 [24]
For calculating $s$-state exciton energies by perturbation theory formulas (2), we need the following matrix elements:
$h_{j}$ $=$ $\langle j|\hat{H}|j\rangle=\frac{\omega}{4}(2j+1)$
$\quad-\frac{2}{\omega\alpha}\left[(2j+1)u_{jj}+ju_{j,j-1}+(j+1)u_{j,j+1}\right],$
$r_{j}$ $=$ $\langle j|\hat{R}|j\rangle=\frac{1}{\omega}(2j+1),$ (13)
$v_{jk}$ $=$ $\langle j|\hat{H}|k\rangle\,\,(j\neq k)=\frac{\omega}{4}\left(k\delta_{j,k-1}+j\delta_{j,k+1}\right)$
$\quad-\frac{2}{\omega\alpha}\left[(2k+1)u_{j,k}+ku_{j,k-1}+(k+1)u_{j,k+1}\right],$
$v^{\text{\tiny R}}_{jk}$ $=$ $\langle j|\hat{R}|k\rangle=\frac{1}{\omega}\left(k\delta_{j,k-1}+j\delta_{j,k+1}\right),$ (14)
where
$u_{jk}(\lambda)$ $=$ $\sum_{s=0}^{\text{min}(k,j)}\sum_{t=0}^{j+k-2s}(-1)^{j+k+t}{{j+k-2s}\choose{t}}$ (15)
×($\qquad\qquad\times{{j}\choose{s}}{{k}\choose{s}}J_{2s+t+1}(\lambda).$
In Eq. (15)), we use the definite integrals
$J_{p}(\lambda)=\int\limits_{0}^{+\infty}\frac{dq}{(1+q)^{p}\sqrt{q^{2}+\lambda^{2}}}$ (16)
with $p\geq 1$ and $\lambda=1/\omega\alpha>0$. These integrals can be calculated numerically or by using the following iterative formulas
$J_{p}=\frac{(2p-3)J_{p-1}-(p-2)J_{p-2}+\lambda}{(\lambda^{2}+1)(p-1)}$ (17)
with $p\geq 2$, where $J_{1}(\lambda)$ has the explicit formula:
$J_{1}(\lambda)=\frac{\ln{\left(\lambda+\sqrt{\lambda^{2}+1}\right)}+\ln{\left(1+\sqrt{\lambda^{2}+1}\right)}-\ln(\lambda)}{\sqrt{\lambda^{2}+1}}.$ (18)
Noting that althought $J_{0}(\lambda)$ is disvergent, relation (17) is still valid for $p=2$ by considering the limit
${\textrm{lim}}_{p\rightarrow 0}\;pJ_{p}(\lambda)=1.$
Function $f_{n}(\lambda)$ in Eq. (3) is calculated by
$f_{n}(\lambda)=u_{n-1,n-1}+\frac{(n-1)u_{n-1,n-2}+nu_{n-1,n}}{2n-1}.$

### v3

@@ Abstract
We have developed an analytical expression for the $s$-state exciton energies in monolayer transition-metal dichalcogenides (TMDCs): $E_{ns}=-{\text{Ry}}^{*}\times P_{n}/{(n-0.5+0.479\,r^{*}_{0}/\kappa)^{2}}$, $n=1,2,...$. Here, $r^{*}_{0}$ and $\kappa$ represent the dimensionless screening length and dielectric constant of the surrounding medium, respectively. $\text{Ry}^{*}$ is an effective Rydberg energy scaled by the dielectric constant and exciton reduced mass, and $P_{n}(r^{*}_{0}/\kappa)$ is a function of variables $n$ and $r^{*}_{0}/\kappa$. Its values are around 1.0, so it is considered a term that corrects the Rydberg energy. Despite the simple form, the suggested formula provides exciton energies with high precision compared to the exact numerical solutions and accurately describes recent experimental data for a large class of TMDC materials, including WSe2, WS2, MoSe2, MoS2, and MoTe2. To achieve these results, we have developed a regulated perturbation theory by combining the conventional perturbation method with several elements of the Feranchuk-Komarov operator method. This includes the Levi-Civita transformation, the algebraic calculation technique using the annihilation and creation operators, and the introduction of a free parameter to optimize the convergence rate of the perturbation series. This universal form of exciton energies could be beneficial in various physical analyses, including retrieval of the material parameters such as reduced exciton mass and screening length from the available measured exciton energies.
@@
Introduction – In the past ten years, energy levels of excitons in different monolayer materials have been measured and studied experimentally and theoretically [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Pioneer studies [1, 2] suggest that, unlike 3D bulk semiconductors, in 2D atomically thin films like monolayer transition-metal dichalcogenides (TMDCs), electrons and holes coupled to each other by what is known as Rytova-Keldysh interaction [19, 20, 21, 22] when they form neutral excitons. This interaction demonstrates a short-range logarithmic potential $\ln(r)$, but becomes a conventional Coulomb interaction $1/r$ in the long range [20, 22]. As a result, the energies of low-lying exciton states (where $n<5$) do not follow the Rydberg series of a two-dimensional hydrogen atom. Therefore, developing an accurate theory of exciton binding energies in 2D semiconductors is essential to benchmark experimental results.
Recently, various methods have been developed to determine exciton binding energies when the reduced mass, screening length, and dielectric constant of the surrounding medium are known [12, 14, 15, 16]. Our packages [14, 15, 16], which are based on the Feranchuk-Komarov operator method [23, 24], perform exceptionally well in calculating exciton binding energies in monolayer TMDCs, even in high magnetic fields. On the other hand, theoretical efforts have been made to derive analytical formulas for exciton binding energies in 2D semiconductors [5, 6, 7]. These formulas result from modeling the electron-hole interaction potential in 2D semiconductors to make the associated Wannier Schrödinger equation analytically solvable. This approach leads to easily applicable formulas that are generalized from the Rydberg series with reasonable accuracy. However, the accuracy of these formulas is, of course, limited by the accuracy of their modified electron-hole potentials. Therefore, a comprehensive theory is required to analytically describe exciton energies.
In this Letter, we present a highly accurate analytical expression for the energies of $s$-state excitons using a new approach called regulated perturbation theory (RPT). The formula we obtained for exciton energies is simple but universally applicable to a wide range of TMDC monolayers. This formula is precise compared to the exact numerical solutions and describes the experimental data that is currently available. The success of this approach is attributed to its advancements. Unlike conventional perturbation theory, the RPT approach incorporates the Levi-Civita transformation, which enables an algebraic calculation technique and introduces a free parameter to control the convergence rate of the perturbation correction series. Furthermore, it eliminates the need to modify the electron-hole potential and accurately computes all matrix elements associated with the harmonic oscillator basis set.
@@
where $\alpha=r_{0}/\kappa a_{0}^{*}$ with $r_{0}$ a screening length.
@@
is excluded from the Coulomb singularity, making solving it more convenient. Notably, it allows the solving process to use the harmonic oscillator basis set $|n,m(\omega)\rangle$ with running indices $n=0,1,2,\cdots$ and $m=0,\pm 1,\pm 2,\cdots,\pm n$ (the principal and magnetic quantum numbers). The harmonic frequency $\omega$ can be used as a free parameter to regulate the convergence rate of the perturbation series so that the desired precision is granted. Moreover, the equation in $(u,v)$ space and the harmonic basis set can be presented via the creation and annihilation operators: $\hat{a}^{+},\hat{a}$ and $\hat{b}^{+},\hat{b}$. All matrix elements necessary for solving Eq. (3), such as $h_{k}=\langle n|\hat{H}|n\rangle$, $r_{k}=\langle n|\hat{R}|n\rangle$, $v_{jk}=\langle j|\hat{H}|k\rangle$, and $v^{\text{\tiny{\it R}}}_{jk}=\langle j|\hat{R}|k\rangle$ $(j\neq k)$, can then be calculated algebraically using the commutation relations $[\hat{a},\hat{a}^{+}]=1$ and $[\hat{b},\hat{b}^{+}]=1$. Here, we consider only $s$-states where $m=0$, so the basis set can be shortly written as $|n\rangle\equiv|n,m=0\rangle$. This algebraic approach obtained analytical forms of these matrix elements. We provide them in Supplement [26] for use in the present study.
@@
Different from Ref. [16], where Eq. (3) was treated numerically with highly accurate exciton energies, we solve this equation analytically by the perturbation method, expanding wave functions and energies by the perturbation parameter $\beta$ as $\psi=\psi^{(0)}+\beta\Delta\psi^{(1)}+\beta^{2}\Delta\psi^{(2)}+\cdots$, $E=E^{(0)}+\beta\Delta E^{(1)}+\beta^{2}\Delta E^{(2)}+\cdots$. Plugging these expansions into Eq. (3), where the operators are separated as $\hat{H}=\hat{H}_{0}+\beta\,\hat{V}$ and $\hat{R}=\hat{R}_{0}+\beta\hat{V}_{R}$, we obtain formulas for the wave function $\psi^{(0)}$ and energy $E^{(0)}$ in the zero-order approximation and their corrections $\Delta\psi^{(s)}$, $\Delta E^{(s)}$ in any approximation order $s=1,2,3,\cdots$. Energies in the $s$-order approximation are calculated by equation $E^{(s)}=E^{(0)}+\Delta E^{(1)}+\Delta E^{(2)}+\cdots+\Delta E^{(s)}$. The present study is interested in exciton energies only to the third-order approximation, whose formulas are provided in Supplement [26].
@@
Table 1 shows exciton energies in monolayer WSe2 in the zero-, second-, and third-order approximations for $1s$ to $5s$ states, calculated by the regulated perturbation theory and compared with the numerically exact solutions [16]. The free parameter $\omega$ is chosen by the variational condition $\partial E^{(0)}/\partial\omega=0$ for these energies. As shown in Table 1, the energy precision is very high, within 1.0 meV for the second-order approximation, while less than 0.2 meV for the third-order approximation. We note that the current limit of the experimental measurement is about 1.0 meV.
@@
Table 1: Exciton energies (meV) in monolayer WSe2 encapsulated by hBN slabs, calculated by the regulated perturbation theory and compared to the exact numerical solutions [16].
$E^{(0)}$ $E^{(2)}$ $E^{(3)}$ $E_{\text{num}}$ [16]
@@
We have also verified the accuracy of the RPT numerically for some other TMDC monolayers: WS2, MoSe2, and MoS2. It shows the same high precision, less than $0.5$ meV, in the third-order approximation, which is not provided here but is available in Supplement [26]. All material parameters used in our calculations are provided in Table 2.
[…]
$\kappa$ $\mu$ ($m_{e}$) $r_{0}$ (nm) Ry (meV) $a^{*}_{0}$ (nm) Refs.
WSe2 4.34 0.19 4.21 137.24 1.209 [16]
WS2 4.16 0.175 3.76 137.59 1.258 [16]
MoSe2 4.40 0.35 3.90 245.97 0.665 [10]
MoS2 4.45 0.275 3.40 188.94 0.856 [10]
MoTe2 4.30 0.37 6.10 272.26 0.615 [26]
Analytical exciton energies – As shown above, the regulated perturbation theory for the corrections of up to the third-order approximation give extremely accurate exciton energies with precision within 0.5 meV compared to the numerically exact solutions. Even for the energies in the second-order approximation, the precision is about 1.0 meV, a limit for experimental measurement. Therefore, we consider them for analytical exciton energies. Indeed, by substituting the explicit forms of matrix elements $h_{k}$, $r_{k}$, $v_{jk}$, and $v^{\text{\tiny{\it R}}}_{jk}$ into RPT equations, we get particular exciton energies in the zero-order approximation as
$E^{(0)}_{ns}=\frac{\kappa^{2}}{4{r_{0}^{*}}^{2}}\frac{1}{\lambda^{2}}-\frac{2\kappa}{r_{0}^{*}}f_{n}(\lambda).$ (4)
Here, $r_{0}^{*}=r_{0}/a_{0}^{*}$ is defined as a dimensionless screening length; $\lambda$ is a variational parameter defined from the free parameter $\omega$ by $\lambda=\kappa/\omega r_{0}^{*}$; $f_{n}(\lambda)$ is a function dependent on $\lambda$ only, whose explicit form is given in Supplement [26]. For convenience, the principle quantum number is defined from now on as $n=1,2,3,\dots$ consistent with the well-known Rydberg series of a hydrogen atom.
@@
$\frac{\partial E^{(0)}_{ns}}{\partial\lambda}=0.$ (5)
We must solve this equation to get $\lambda$ and plug it into Eq. (4) to get the needed analytical energies. We demonstrate the solving process for $ns$ states ($n\leq 5$) in Supplement [26] and provide only the main idea and final results in the main text.
Because of the complexity of function $f_{n}(\lambda)$, analytically solving Eq. (5) is not trivial, requiring the Taylor expansion of this function. From another side, numerically solving this equation for specific cases of monolayers WSe2, WS2, MoSe2, and MoS2 encapsulated by hBN slabs with the material parameters given in Table 2, we get particular solutions for $1s$ state: $\lambda=0.803,0.863,0.582,$ and $0.750$, corresponding to each material. These results suggest we expand function $f_{1}(\lambda)$ around the average value $\lambda_{0}=0.75$. This expansion leads Eq. (5) to a polynomial equation concerning an unknown $\delta\lambda=\lambda-0.75$, making it easy to get approximate analytical solutions consistent with the numerical ones. Analogically, we get $\lambda$ for other $ns$ states by expanding functions $f_{n}(\lambda)$ around the points $\lambda_{0}=(n+0.5)/2$ and then solving Eq. (5). We combine all the solutions in one generalized formula as $\lambda=\lambda_{0}+\delta\lambda$ with
$\delta\lambda=\lambda-(n+0.5)/2=\left(0.47\,{\kappa}/{r_{0}^{*}}-0.525\right)n\,.$ (6)
We note that the actual solutions of Eq. (5) are not precisely as given by formula (6); however, a slight modification has been made to get a general formula for all states $ns$. This approach is reasonable because exact energies should not depend on the free parameter, and its modification around the extremum point does not change the solutions much. Reference [24] provides a detailed discussion of how to choose such a free parameter.
@@
$E_{ns}$ $=$ $-\frac{{\text{Ry}}^{*}\times P_{n}}{(n-0.5+0.479\,r_{0}^{*}/\kappa)^{2}}.$ (7)
@@
with $\Delta=(0.47\,{\kappa}/{r_{0}^{*}}-0.525)$, coefficients $a_{n}=4.3/n-{6.6}/{n^{2}}+{3.2}/{n^{3}}$ and $b_{n}=5.3/n-{7.6}/{n^{2}}+{4.0}/{n^{3}}.$ $P_{n}$ is a function of variable $r_{0}^{*}/\kappa$ only with values around 1.0 and $P_{n}\rightarrow 1$ when $n\rightarrow+\infty$, so we can interpret it effective screening charge correction.
@@
Validation with the available experimental data – To obtain analytical energies (7), we expand the equation around the specific points of $\lambda$ with the condition $\lambda\gg\delta\lambda$, which defines the boundary of our theory, the working range of $r_{0}^{*}/\kappa$ as
$0.46\ll r_{0}^{*}/\kappa\ll 18.8$ (9)
By definition, $r^{*}_{0}/\kappa=\mu/\kappa^{2}\times r_{0}/a_{0}$, $a_{0}=0.0529$ nm - Bohr radius, so boundary (9) leads to a working range for $\kappa$, in particular for WSe2, $0.89\ll\kappa\ll 5.73$.
@@
Furthermore, we verify formula (7) by comparing it with experimental data for monolayers WSe2 [9], WS2, MoSe2, MoS2, and MoTe2 [10]. The comparison in Fig. S-1 demonstrates a high consistency between the experiments and theory. For MoTe2, material parameters are given in [10] with a large uncertainty. However, we choose their values (in Table 2) so that our theory best fits the experimental data.
Experimental data of Liu et al. [8] for monolayer WSe2 also support our theory: $\Delta E_{2s-1s}=131$(130.3) meV, $\Delta E_{3s-2s}=21$ (21.7) meV, $\Delta E_{4s-2s}=30$ (29.0) meV, where the values in parenthesis are theoretical from Eq. (7).
@@
Figure 1: Exciton binding energies $E_{b}$ for the $ns$ states in monolayers WSe2, WS2, MoSe2, MoS2, and MoTe2 calculated by formula (7) (horizontal axis) compared with the experimental data [10, 9] (vertical axis). Material parameters used for calculations are provided in Table 2.
@@
Conclusion – A novel approach based on the perturbation theory has been developed, incorporating some elements of the FK operator method to achieve the energies of $s$-state excitons in monolayer transition-metal dichalcogenides. The third-order approximation of perturbation theory presents almost exact solutions, while the second-order approximation gives exciton energies with high precision, within 1.0 meV compared to the exact numerical solutions. This accuracy suggests we use them for analytical exciton energies. Although simple and universal for $s$-states, the constructed analytical formula describes experimental exciton energies for a wide range of monolayer TMDCs. This general approach could be applied to other 2D semiconduction materials.
@@
Acknowledgments – This work was funded by Vietnam Ministry of Education and Training under grant number B2022-SPS-09-VL and carried out by the high-performance cluster at Ho Chi Minh City University of Education, Vietnam.
Author contribution – V.-H. Le conceptualized the work; V.-H. Le, H.T. Dinh, Dai-Nam Le, and N.-H. Phan performed analytical formulation, N.-H. Phan, Duy-Nhat Ly, N.-Q. Nguyen, and P.-T. Doan carried out numerical calculations; V.-H. Le, Dai-Nam Le, H.T. Dinh, and N.-T. D. Hoang analyzed data; V.-H. Le, Dai-Nam Le, and N.-H. Phan wrote and edited the manuscript.